• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 15:56:05 on March 21, 2011

    Entry number 352735

    keyword=MCC update

    The power supply was brought up and the problem seems to have disappeared, we might have beam around 8 PM tonight, but they will decide after a few hours of testing/checking. So I will put the swing shift on stand by only the target operator has to stay but people should stay around the lab for now until we have a better estimate. People on shift should leave a number where they can be reached.



    A copy of this log entry has been emailed to: norum